OpenStudy (anonymous):

If the geometric mean of a and 176 is 44, find the value of a.???? explain

OpenStudy (amistre64):

geo means are when you multiply it all together and take the nth root

OpenStudy (amistre64):

sqrt(176a) = 44 ; solve for a

OpenStudy (amistre64):

do you see how we could start to solve for "a" ?

OpenStudy (anonymous):

We have 2 values. (176 + a)/2 = 44 So we need to do the algebra on it and solve for a. I am just getting off a train. Will check back in a while and see if you still need help. Best of luck! http://www.tutorsean.net

OpenStudy (anonymous):

Another way to look at geo mean is to use the ratio equation\[\frac{x}{a}=\frac{a}{y}\] where a is the geo mena of x and y.

OpenStudy (amistre64):

sorry sean; thats arithmetic mean; solve for the geometric mean

OpenStudy (amistre64):

ive never tried the ratio b4 :)

OpenStudy (anonymous):

so in this case we have\[\frac{a}{44}=\frac{44}{176}\]

OpenStudy (anonymous):

oh thank u all very helpful

OpenStudy (amistre64):

how would the ratio work for say, 3 terms?

OpenStudy (anonymous):

As far as i can tell, it doesn't. But i remember learning geo mean from "Donald in Mathemagic-land" years ago. It applies to the Golden Ratio and they were talking about how it relates to the dimensions of the rectangle.